OpenStudy (anonymous):

For the following graphs, state the intervals of x for which the functions are increasing and those for which they are decreasing. Sketch the graphs- (iv) y= 3x^2 - (1/x^2), (x does not equal 0).

OpenStudy (anonymous):

... is this calc or algebra?

OpenStudy (anonymous):

Calculus again.

OpenStudy (anonymous):

oh ok. so take the derivatives. if first derivative > 0 , increasing, <0 decreasing, =0 it's flat / not changing.

OpenStudy (anonymous):

That's my problem. the derivative I've got here is x^4+(1/3)

OpenStudy (anonymous):

x^4 + (1/3) > 0. How do you solve that?

OpenStudy (anonymous):

In these questions, you don't use "i".

OpenStudy (anonymous):

well i can't tell ... use the equation editor please

OpenStudy (anonymous):

\[\frac{ d }{ dx } x^4 + \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

no i mean the original equation. sorry.

OpenStudy (anonymous):

Oops, wait that's the wrong derivative sorry. That was simplified when I made the derivative less than zero. \[y= 3x^2 - \frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

\[\frac{ d }{ dx } \frac{ 6x^4 + 2 }{ x^3 }\]

OpenStudy (anonymous):

huh. oh sorry yes?

OpenStudy (anonymous):

by the way the derivative is \[12x-\frac{ 4 }{ x^3 }\]

OpenStudy (anonymous):

so personally i don't like that form so let's make it into a combined form ex. a/b where a and b are expressions

OpenStudy (anonymous):

Wait. @RONNCC I combined the equation before differentiating?

OpenStudy (anonymous):

Are you meant to differentiate then combine?

OpenStudy (amoodarya):

look at image if x>0 f increase if x<0 f dec

OpenStudy (anonymous):

f increases when x<0, that's what the answers say.

OpenStudy (anonymous):

I agree with your diagram but why does f increase when x<0. That's what the answers say.

OpenStudy (anonymous):

well the function is increasing whenever its derivative is > 0, and decreasing whenever its derivative is < 0

OpenStudy (anonymous):

The answers say decreasing when x< 0 increasing when x < 0.

OpenStudy (anonymous):

if f(x) = 3x^2 - 1/x^2 then f'(x) = \[\frac{ 2 }{ x^{3} } + 6x\] or combined \[\frac{ 6x^{4} + 2 }{ x^{3} }\]

OpenStudy (anonymous):

we know the numerator is always positive so we must look at the denominator. whenever the denominator is positive we have +/+ which is positive so if x > 0, the derivative is positive if the denominator is negative we have +/- which is negative so if x < 0 the derivative is negative

OpenStudy (anonymous):

"The answers say decreasing when x< 0 increasing when x < 0" what does that mean? it cant be decreasing when x < 0 and increasing when x < 0. that is a typo

OpenStudy (anonymous):

That's what it says.

OpenStudy (anonymous):

it contradicted itself then

OpenStudy (anonymous):

Is it meant to be increasing when x>0?

OpenStudy (anonymous):

probably lol

OpenStudy (anonymous):

thats like saying a number is positive, but also negative

OpenStudy (anonymous):

I think it means the derivative is negative when x<0 and the derivative is positive when x<0

OpenStudy (anonymous):

i think it means derivative is negative if x < 0 derivative is positive if x > 0 because that's what the first derivative shows

OpenStudy (anonymous):

ah okay